# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

### v1

@@ 3.2 Reverse process
Secondly, we reference the reverse SDE (see eq. 15) and incorporating the prediction of $\mathbf{x}$:
$d\mathbf{z}_{t}=$ $\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}$ (16)
$\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)},$ (17)
@@
However, to fully specify the reverse process, it is necessary to define a prior distribution $p(\mathbf{z}_{1})$ and a reconstruction distribution $p(\mathbf{x}|\mathbf{z}_{0})$. In all of our experiments, we set the prior $p(\mathbf{z}_{1})$ to be a unit Gaussian distribution $\mathcal{N}(\mathbf{z}_{1};0,I)$ and let $p(\mathbf{x}|\mathbf{z}_{0})$ be a Gaussian distribution with a small variance $\mathcal{N}(\mathbf{x};\mathbf{z}_{0},\delta^{2}I)$, where $\delta^{2}=10^{-4}$.
This particular parameterization of the reverse process is not the only possibility. However, it is a convenient choice as it allows for the definition of the reverse process simply through the prediction of $\mathbf{x}$, akin to conventional diffusion models (Ho et al., 2020).
3.3 Optimization
We propose to optimize the forward and reverse processes of NFDM jointly, minimizing the following objective:
$\mathcal{L}={\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}},$ (18)
${\mathcal{L}_{\mathrm{rec}}}=\underset{q_{\varphi}(\mathbf{x},\mathbf{z}_{0})}{\mathbb{E}}\left[-\log p(\mathbf{x}|\mathbf{z}_{0})\right],$ (19)
${\mathcal{L}_{\mathrm{diff}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right],$ (20)
${\mathcal{L}_{\mathrm{prior}}}=\underset{q(\mathbf{x})}{\mathbb{E}}\left[{\mathrm{D}_{\mathrm{KL}}}\big{(}q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\|p(\mathbf{z}_{1})\big{)}\right].$ (21)
As we demonstrate in Section A.1 this objectives provide a variational bound on the model’s likelihood $p_{\theta,\varphi}(\mathbf{x})$.
The objective $\mathcal{L}$ shares similarities with the standard diffusion model objective (Ho et al., 2020), though the individual terms differ. Moreover, if the forward process is parameterized by $\varphi$, we need to optimize the objective (eq. 18) with respect to these parameters. The objective $\mathcal{L}$ also exhibits strong connections with both the Flow Matching (Lipman et al., 2022) and Score Matching (Vincent, 2011) objectives. We explore these connections in Section B.2, where we also discuss the role of $g_{\varphi}$.
A key characteristic of the NFDM objective is its amenability to training within a simulation-free paradigm, which is critical for efficient optimization. We summarize the training procedure in Algorithm 1.
@@ 3.3 Optimization
0:  $q(x)$, $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$
@@
     $\mathbf{x}\sim q(x)$, $t\sim u(t)$, $\{\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{1}\}\sim q(\cdot|\mathbf{x})$
     ${\mathcal{L}_{\mathrm{rec}}}=-\log p(\mathbf{x}|\mathbf{z}_{0})$
     ${\mathcal{L}_{\mathrm{diff}}}=\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}$
     ${\mathcal{L}_{\mathrm{prior}}}=\log q_{\varphi}(\mathbf{z}_{1}|x)-\log p(\mathbf{z}_{1})$
     Gradient step on $\theta$ and $\varphi$ w.r.t. ${\mathcal{L}_{\mathrm{rec}}}+{\mathcal{L}_{\mathrm{diff}}}+{\mathcal{L}_{\mathrm{prior}}}$
@@
0:  $F_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$, $T$ – number of steps
@@
3.4 Sampling
NFDM offers several methods for sampling from the trained reverse process. The first method is stochastic sampling, as defined in Section 3.2. This process involves the following steps: 1) sample $\mathbf{z}_{1}\sim p(\mathbf{z}_{1})$; 2) numerically simulate the SDE (eq. 16) to obtain $\mathbf{z}_{0}$; 3) sample $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$. This procedure is summarized in Algorithm 2.
@@ 3.4 Sampling
Moreover, during the sampling process, we can adjust the level of stochasticity by modifying $g_{\varphi}(t)$ (eq. 12). It is important to note that changes to $g_{\varphi}(t)$ also influence $\tilde{f}^{F}_{\varphi}$ (eq. 12) and $\hat{f}_{\theta,\varphi}$ (eq. 17). In the extreme case where $g_{\varphi}(t)\equiv 0$, the reverse process becomes deterministic, allowing us to utilize off-the-shelf numerical ODE solvers and to estimate densities (Chen et al., 2018a; Grathwohl et al., 2018).
3.5 Parameterization
The parameterization of the variance function $g_{\varphi}(t)$ (eq. 12) and the data point predictor $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (eq. 17) is straightforward. In our experiments, we parameterize them directly using neural networks. However, the parameterization of the transformation $F_{\varphi}$ (eq. 9) is more complex. NFDM necessitates that $F_{\varphi}$ be differentiable and invertible with respect to $\varepsilon$, as well as provide access to the logarithm of the determinant of the transformation $\log|J_{F}^{-1}|$ (eq. 14). Therefore, we use the following parameterization for $F_{\varphi}$:
$\sigma_{\varphi}(\mathbf{x},t)$ $=\delta+t\bar{\sigma}_{\varphi}\big{(}(1-t)\mathbf{x},t\big{)}/\bar{\sigma}_{\varphi}(\mathbf{0},1).$ (24)
In this context, $\bar{\mu}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}$ and $\bar{\sigma}_{\varphi}:\mathbb{R}^{D}\times[0,1]\mapsto\mathbb{R}^{D}_{+}$, with $\delta$ being a constant. In our experiments, we set $\delta^{2}=10^{-4}$. If $\bar{\mu}_{\varphi}$ and $\bar{\sigma}_{\varphi}$ are differentiable, this parameterization meets the requirements on $F_{\varphi}$ and renders $q(\mathbf{z}_{t}|\mathbf{x})$ a conditional Gaussian distribution with parameterized mean and diagonal covariance matrix.
Furthermore, this reparameterization ensures that $q(\mathbf{z}_{0}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{0};\mathbf{x},\delta^{2}I)$ and $q(\mathbf{z}_{1}|\mathbf{x})=\mathcal{N}(\mathbf{z}_{1};0,I)$. These restrictions are not necessary for $F_{\varphi}$, however, it eliminates the need to optimize the reconstruction loss ${\mathcal{L}_{\mathrm{rec}}}$ (eq. 19) and the prior loss ${\mathcal{L}_{\mathrm{prior}}}$ (eq. 21), as they are independent of the parameters $\varphi$ and $\theta$.
The chosen parameterization of $F_{\varphi}$, $g_{\varphi}(t)$, and $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ is not the only possible approach and may not be optimal. For instance, it is possible to parameterize $F_{\varphi}$ as a flow-based model (Dinh et al., 2016; Kingma & Dhariwal, 2018), enabling a broader range of distributions than just conditional Gaussians. However, for the purpose of this paper, we maintain a simple setup and leave the exploration of the most effective parameterization for future research.
We discuss the parameterization in more detail in Section C.2.
@@ 3.5 Parameterization
Table 1: Comparison of NFDM results with baselines on density estimation tasks. We present results in terms of BPD, lower is better. NFDM achieves state-of-the-art results accross the three benchmark tasks.
@@
DDPM (Ho et al., 2020) $3.69$
Score SDE (Song et al., 2020b) $2.99$
Improved DDPM (Nichol & Dhariwal, 2021) $2.94$ $3.54$
VDM (Kingma et al., 2021) $2.65$ $3.72$ $3.40$
Score Flow (Song et al., 2021) $2.83$ $3.76$
Flow Matching (Lipman et al., 2022) $2.99$ $3.53$ $3.31$
Stochastic Interp. (Albergo & Vanden-Eijnden, 2022) $2.99$ $3.48$
i-DODE (Zheng et al., 2023b) $2.56$ $3.43$
MuLAN (Sahoo et al., 2023) $2.55$ $3.67$
NFDM (this paper) $\mathbf{2.49}$ $\mathbf{3.36}$ $\mathbf{3.20}$
4 Restricted NFDM
We introduce NFDM as a powerful framework that enables both pre-definition and learning of the forward process, potentially simplifying the task for the reverse process. However, there are in general an infinite number of forward and reverse processes that correspond to each other. So how to choose between them? The standard NFDM, as described in Section 3, learns one such pair of processes. However, the flexibility of NFDM opens up the possibility of choosing specific pairs of processes depending on the task.
Suppose our objective is not just to learn any generative process for the data, but to learn one endowed with specific beneficial properties. For instance, one such property is straight generative ODE trajectories. Learning a generative process with straight ODE trajectories can be highly beneficial, as it enables the generation of samples with far fewer steps in the ODE solver. This can lead to significant computational savings.
To achieve this, we must introduce some restrictions on the reverse process. One approach is to directly encode this in the definition of the reverse process. Alternatively, we can introduce penalties on the curvature of its trajectories. In both cases learnable forward process would then adapt to align with the reverse process. As a result, we would obtain a generative process that not only corresponds to the forward process, ensuring accurate data generation, but also features the desired property of straight trajectories. We discuss NFDM with restrictions in more details in Section A.2.
Importantly, conventional diffusion models are incapable of handling such constraints or penalization strategies. In diffusion models with a fixed forward process, the target for the reverse process is predetermined and the corresponding ODE trajectories are highly curved. Hence, imposing constraints on the reverse process, such as trajectory straightness, would lead to a mismatch with the forward process and, consequently, an inability to generate samples with high data fidelity.
4.1 NFDM with optimal transport
In this section, we discuss curvature penalization on the ODE trajectories of the reverse process in NFDM. We refer to this variant as NFDM-OT [^1] OT stands for optimal transport. This designation is used for convenience. While straight trajectories are a necessary condition for dynamic optimal transport, they are not sufficient..
We propose learning a model with an additional curvxature penalty as suggested by Kelly et al. (2020):
${\mathcal{L}_{\mathrm{OT}}}$ $=\mathcal{L}+\lambda{\mathcal{L}_{\mathrm{crv}}},\quad\textrm{where}$ (25)
${\mathcal{L}_{\mathrm{crv}}}=\underset{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}{\mathbb{E}}$ $\left\|\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}\right\|_{2}^{2}.$ (26)
${\mathcal{L}_{\mathrm{crv}}}$ is an additional curvature loss that penalizes the second time derivative of the generative ODE trajectories. $\mathcal{L}$ is estimated as in section 3.3, whereas when calculating ${\mathcal{L}_{\mathrm{crv}}}$ we set $g_{\varphi}(t)\equiv 0$. ${\mathcal{L}_{\mathrm{crv}}}\equiv 0$ ensures that the generative trajectories are straight lines. Although we parameterize $\hat{f}_{\theta,\varphi}$ (eq. 17) through $\tilde{f}^{B}_{\varphi}$ (eq. 15), ${\mathcal{L}_{\mathrm{crv}}}$ does not penalize the curvature of the conditional forward trajectories. The penalty requires $F_{\varphi}$ to be twice differentiable with respect to time. The specifics of the curvature loss are elaborated upon in more detail in Section A.2.
Even though the parameterization, as discussed in section 3.5, is simple and does not guarantee reaching the zero curvature, we find that in practice it is flexible enough to successfully learn straight line generative trajectories. In our experiments, we set $\lambda=10^{-2}$. Empirical evidence suggests that higher values of $\lambda$ lead to slower convergence of the model. We provide empirical results for NFDM-OT in Section 5.2.

### v2

@@ 3.2 Reverse (Generative) Process
Secondly, leveraging this reverse SDE (see eq. 10) we incorporate a prediction of $\mathbf{x}$:
$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)\big{)},$ (11)
@@
To fully specify the reverse process, it is also necessary to define a prior distribution $p(\mathbf{z}_{1})$ and a reconstruction distribution $p(\mathbf{x}|\mathbf{z}_{0})$. In all of our experiments, we set the prior $p(\mathbf{z}_{1})$ to be a unit Gaussian distribution $\mathcal{N}(\mathbf{z}_{1};0,I)$ and let $p(\mathbf{x}|\mathbf{z}_{0})$ be a Gaussian distribution with a small variance $\mathcal{N}(\mathbf{x};\mathbf{z}_{0},\delta^{2}I)$, where $\delta^{2}=10^{-4}$.
The above parameterization of the reverse process is not the only possibility. However, it is a convenient choice as it allows for the definition of the reverse process simply through the prediction of $\mathbf{x}$, similar to conventional diffusion models [19]. We leave exploration of alternate parameterizations for future work.
3.3 Optimization and Sampling
With $F_{\varphi}$ in eq. 7 parameterized such that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})\approx p(\mathbf{z}_{1})$, we propose to optimize the forward and reverse processes of NFDM jointly, minimizing the following objective:
$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right].$ (12)
Since the forward process is parameterized by $\varphi$, we need to optimize the objective (eq. 12) with respect to both $\varphi$ and $\theta$ jointly. We discuss parameterization of the forward process in Section B.5.
As we demonstrate in Section A.1 the objective $\mathcal{L}$ shares similarities with the standard diffusion model objective [19], and it provides a variational bound on the model’s log-likelihood $\log p_{\theta,\varphi}(\mathbf{x})$. The objective $\mathcal{L}$ also exhibits strong connections with both the Flow Matching [32] and Score Matching [67] objectives. We explore these connections in Section E.4, where we also discuss the role of $g_{\varphi}$ in detail. It is important to note that despite the parameterization of the reverse process through the prediction of $\mathbf{x}$, the objective $\mathcal{L}$ optimizes the generative dynamics and does not necessarily lead to accurate predictions of $\mathbf{x}$.
A key characteristic of the NFDM objective is its compatibility with the simulation-free paradigm, which is critical for efficient optimization. We summarize the training procedure in Algorithm 1.
@@ 3.3 Optimization and Sampling
$q(\mathbf{x})$, $F_{\varphi}$, $g_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$
@@
     $\mathbf{x}\sim q(\mathbf{x})$, $t\sim u(t)$
     $\mathbf{z}_{t}\sim q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$
     $\mathcal{L}=\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}$
     Gradient step on $\theta$ and $\varphi$ w.r.t. $\mathcal{L}$
@@
$F_{\varphi}$, $g_{\varphi}$, $\hat{\mathbf{x}}_{\theta}$, $T$ – number of steps
@@
To sample from the trained reverse process we can simulate the SDE, as defined in Section 3.2. This procedure is summarized in Algorithm 2. Additionally, during the sampling process, we can adjust the level of stochasticity by modifying $g_{\varphi}(t)$ (eq. 9). It is important to note that changes to $g_{\varphi}(t)$ also influence $\tilde{f}^{F}_{\varphi}$ (eq. 9) and $\hat{f}_{\theta,\varphi}$ (eq. 11). In the extreme case where $g_{\varphi}(t)\equiv 0$, the reverse process becomes deterministic, allowing us to utilize off-the-shelf numerical ODE solvers and to estimate densities [6, 16]. We provide an extended sampling discussion in Section B.4.
4 Neural Flow Bridge Models
In this section we discuss a simple modification to the NFDM framework that enables us to learn bridges between two data distributions, $q(\mathbf{x}_{0})$ and $q(\mathbf{x}_{1})$, a modification we refer to as Neural Flow Bridge Models (NFBM).
[…]
$\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x}_{0},\mathbf{x}_{1}),$ (13)
Similar to the discussion in Section 3.2, for the reverse process of NFBM, we predict the data points $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ from $\mathbf{z}_{t}$ and $t$, and substitute them into the conditional reverse time SDE (eq. 10):
$d\mathbf{z}_{t}=\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)dt+g_{\varphi}(t)d\bar{\mathbf{w}},\quad\textrm{where}\quad\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)=\tilde{f}^{B}_{\varphi}\big{(}\mathbf{z}_{t},t,\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)\big{)}.$ (14)
Here, the function $\hat{\mathbf{x}}^{0,1}_{\theta}(\mathbf{z}_{t},t)$ returns predictions of both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. Alternatively, the data point $\mathbf{x}_{1}$ can be reused as conditioning for intermediate steps of the reverse process instead of predicting both points, a strategy further detailed in Section C.2.
When $F_{\varphi}$ in eq. 13 is parameterized such that $q_{\varphi}(\mathbf{z}_{0}|\mathbf{x}_{0},\mathbf{x}_{1})\approx\delta(\mathbf{x}-\mathbf{z}_{0})$ and $q_{\varphi}(\mathbf{z}_{1}|\mathbf{x}_{0},\mathbf{x}_{1})\approx p(\mathbf{z}_{1}|\mathbf{x}_{1})$ (see Section C.1 for parameterization details of the NFBM forward process), we propose training NFBM by minimizing the following objective:
$\mathcal{L}=\mathbb{E}_{u(t)q(\mathbf{x}_{0},\mathbf{x}_{1})q_{\varphi}(\mathbf{z}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{1})}\left[\frac{1}{2g_{\varphi}^{2}(t)}\big{\|}\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x}_{0},\mathbf{x}_{1})-\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)\big{\|}_{2}^{2}\right].$ (15)
The derivation of this objective is provided in Section A.2. This objective shares key properties with the NFDM objective, it provides a variational bound on the model’s log-likelihood $\log p_{\theta,\varphi}(\mathbf{x}_{0})$, and is compatible with the simulation-free paradigm. Furthermore, it allows for sampling with various levels of stochasticity (by adjusting $g_{\varphi}(t)$) including deterministic sampling (when $g_{\varphi}(t)\equiv 0$).
Thus, the NFDM framework not only enables the construction of generative models, but also facilitates learning bridge models between two data distributions.
5 Restricted NFDM
We introduce NFDM as a powerful framework that enables learning of the forward process. However, there is in general an infinite number of forward and reverse processes that correspond to each other. In this section we discuss how the flexibility of NFDM (and NFBM) allows learning generative dynamic with user-specified beneficial properties.
Suppose our objective is to learn straight generative ODE trajectories, which can be highly beneficial, as it enables generation with far fewer steps in the ODE solver. One approach is to introduce penalties on the curvature of the reverse process’s trajectories and let the learnable forward process to adapt to align with the reverse process. As a result, we would obtain a generative process that not only corresponds to the forward process, ensuring accurate data generation, but also features the desired property of straight trajectories. We discuss NFDM with restrictions in more details in Section D.1.
We propose learning a model with curvature penalty as suggested by [24]:
${\mathcal{L}_{\mathrm{OT}}}=\mathcal{L}+\lambda{\mathcal{L}_{\mathrm{crv}}},\quad\textrm{where}\quad{\mathcal{L}_{\mathrm{crv}}}=\mathbb{E}_{u(t)q_{\varphi}(\mathbf{x},\mathbf{z}_{t})}\left\|\frac{d\hat{f}_{\theta,\varphi}(\mathbf{z}_{t},t)}{dt}\right\|_{2}^{2}.$ (16)
We refer to this variant as NFDM-OT [^1] OT stands for optimal transport. This designation is used for convenience. While straight trajectories are a necessary condition for dynamic optimal transport, they are not sufficient.. ${\mathcal{L}_{\mathrm{crv}}}$ is an additional curvature loss that penalizes the second time derivative of the generative ODE trajectories. $\mathcal{L}$ is estimated as in Section 3.3, whereas when calculating ${\mathcal{L}_{\mathrm{crv}}}$ we set $g_{\varphi}(t)\equiv 0$. ${\mathcal{L}_{\mathrm{crv}}}\equiv 0$ ensures that the generative trajectories are straight lines. In our experiments, we set $\lambda=10^{-2}$. Empirical evidence suggests that higher values of $\lambda$ lead to slower convergence of the model. The specifics of the curvature loss are elaborated upon in more detail in Section D.2.
We would like to emphasise that the purpose of this section is to provide an example of how the NFDM framework allows learning dynamics with specific properties. Not to propose the best way for learning generative straight-line generative dynamics. In addition, we note that conventional diffusion models are incapable of handling such penalization strategies. In diffusion models with a fixed forward process, the target for the reverse process is predetermined and the corresponding ODE trajectories are highly curved. Hence, imposing constraints on the reverse process, such as trajectory straightness, would lead to a mismatch with the forward process and, consequently, an inability to generate samples with high data fidelity.
6 Experiments
We first showcase results demonstrating that NFDM consistently achieves better likelihood compared to baselines, obtaining state-of-the-art diffusion modeling results on the CIFAR-10 [30] and downsampled ImageNet [11, 66] datasets. Then, we explore the NFDM-OT modification, which penalizes the curvature of the deterministic generative trajectories. The NFDM-OT reduces trajectory curvature, significantly lowering the number of generative steps required for sampling. Finally, we demonstrate the NFBM modification, learning bridges on the AFHQ [8] dataset and several synthetic examples.
We report the NLL in Bits Per Dimension (BPD) and sample quality measured by the Frechet Inception Distance (FID) score [18]. The NLL is calculated by integrating the ODEs using the RK45 solver [15], with all NLL metrics computed on test data. For FID, we provide the average over 50k generated images.
Unless otherwise stated, we parameterized the forward process of NFDM and NFBM such that $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ is a Gaussian with learnable mean and covariance (see details in Sections B.5 and C.1). For a detailed description of parameterizations and other experimental details, please refer to Appendix F.
The primary aim of the experiments with NFDM-OT and NFBM is not to introduce a novel model that surpasses others in few-step generation or learning bridges, but rather to showcase the ability of the NFDM framework to learn generative dynamics with specific properties. The straightness of the trajectories is just one example of such a property. We leave it for future research to explore different parameterizations and modified objectives for NFDM that may yield even better results.
Please refer to section G.3 for generated samples.
6.1 Likelihood Estimation
@@
Table 1: Comparison of NFDM results with baselines on density estimation tasks. We present results in terms of BPD, lower is better. NFDM achieves state-of-the-art results across all three benchmark tasks.
@@
DDPM [19] $3.69$
Score SDE [60] $2.99$
Improved DDPM [38] $2.94$ $3.54$
VDM [27] $2.65$ $3.72$ $3.40$
Score Flow [58] $2.83$ $3.76$
Flow Matching [32] $2.99$ $3.53$ $3.31$
Stochastic Interp. [2] $2.99$ $3.48$
i-DODE [74] $2.56$ $3.43$
NDM [4] $2.70$ $3.55$ $3.35$
MuLAN [48] $2.55$ $3.67$
NFDM (Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$) $2.49$ $3.36$ $\mathbf{3.20}$
NFDM (non-Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$) $\mathbf{2.48}$ $\mathbf{3.34}$ $\mathbf{3.20}$
(a) Score SDE [60]
(b) NFDM-OT (this paper)
Figure 1: Comparison of trajectories between the data distribution (on the left) and the prior distribution (on the right), as learned by conventional diffusion and NFDM-OT.
For the first experiments, in addition to Gaussian we also provide results with non-Gaussian parameterization of the forward process (see Section B.5). Table 1 summarizes the NLL results on the CIFAR-10 and two downsampled ImageNet datasets. Notably, NFDM outperforms diffusion-based baselines on all three datasets with both parameterizations, achieving state-of-the-art performance. The improved performance due to NFDM is a natural progression for the following reasons.
First, it is well-established [58, 74] that diffusion models exhibit improved likelihood estimation when trained with the full ELBO objective. The objective $\mathcal{L}$ (eq. 12) used for training the NFDM is also a variational bound on the likelihood.
Second, diffusion models can be seen as hierarchical VAEs. From this perspective, most baselines in Table 1 resemble VAEs with either fixed or constrained variational distributions. In contrast, the NFDM extends beyond these baselines by providing a more flexible variational distribution. That is true even for the Gaussian parameterization with a mean and covariance that is non-linear in $\mathbf{x}$ and $t$. This flexibility allows the NFDM to better conform to the reverse process, consequently enhancing likelihood estimation.
6.2 Straight Trajectories
We next evaluate NFDM-OT, designed to penalize the deterministic generative trajectory curvature. First, we compare NFDM-OT with a conventional continuous-time diffusion model. Figure 1(a) illustrates deterministic trajectories between a two-dimensional data distribution and a unit Gaussian distribution learnt by a conventional diffusion model [60]. Figure 1(b) depicts trajectories learnt by NFDM-OT. Conventional diffusion, being constrained in its forward process, learns highly curved trajectories, whereas NFDM-OT successfully learns straight generative trajectories as desired. In Section G.1 we provide some additional reult, demonstrating the importance of learnable forward process for NFDM-OT.
Then, we present results of NFDM-OT on image datasets. Table 2 reports the FID scores for $2$, $4$, and $12$ Number of Function Evaluations (NFE) with respect to the function $\hat{f}_{\theta,\varphi}$ (eq. 11). In this experiment, we employ Euler’s method for sampling integration. For the specified NFEs, NFDM-OT demonstrates superior sample quality compared to other approaches with similar NFE values. Specifically, NFDM-OT outperforms approaches specifically designed to minimize the curvature of generative trajectories [43, 31].
Importantly, NFDM-OT is trained with an ELBO-based objective (eq. 12), which is known to yield higher FID scores for diffusion models [58, 74]. In contrast, some of the approaches listed in Table 2 are trained with different objectives, leading to improved FID scores. Even so, for comparable NFE values NFDM-OT still achieves superior results. We provide additional results of NFDM-OT in Section G.2.